At a store, you are interested in purchasing a candy bar and a piece of bubble gum. Together, they cost $1.20. The candy bar costs a dollar more than four times the bubble gum. How much does the piece of bubble gum cost, in cents?
Answer by math_tutor2020(3817)


b = cost of bubble gum
4b+100 = cost of candy bar
Both costs are in cents.

candybar + gum = 120 cents
(4b+100)+b = 120
5b+100 = 120
5b = 120-100
5b = 20
b = 20/5
b = 4
The gum costs 4 cents.

4b+100 = 4*4+100 = 116 cents, aka $1.16, is the cost of the candy bar.
candybar + gum = $1.16 + $0.04 = $1.20 which confirms the answer is correct.


Answer: 4 cents